OpenStudy (anonymous):

Use the formula, A=P(1+rt) A=$1620 P=$1500 t=4 years

OpenStudy (anonymous):

27%?

OpenStudy (jackellyn):

Substitute the values into your equation 1620=1500(1+r(4)) Distribute 1620 = 1500 + 1500(4r) Simplify and solve for r

OpenStudy (anonymous):

It's a little smaller than the answer you've got. Jackellyn's answer has the method down.

OpenStudy (anonymous):

7.4?

OpenStudy (anonymous):

wait hold on i solved wrong

OpenStudy (anonymous):

is it 0.02?

OpenStudy (jackellyn):

0.02 is correct @monroe17

OpenStudy (anonymous):

0.02% or 2%?

OpenStudy (anonymous):

As there are no percents in the original equation, the answer of 0.02 is not a percent. The correct answer would be 0.02. 2% is a little confusing because we might not know what it is 2% of...

OpenStudy (anonymous):

but the rate is in percentage..

OpenStudy (anonymous):

Fair enough, then. You know the problem and the course better than I. If there's no risk of confusion, go for it!

OpenStudy (anonymous):

Having reread the problem - the use of a percent here would be perfectly natural.

OpenStudy (anonymous):

Yeah, I got it correct. Thanks:)